The second harmonic of an open pipe has a frequency double the frequency of the third harmonic of a closed pipe 5 m long. How long is the open pipe?


1
Expert's answer
2021-05-11T03:52:03-0400

Second harmonic of open pipe=v/l

Where v is the velocity, l is the length of open pipe

Third harmonic of closed pipe=3v/4L

Where L is the length of closed pipe=5m

According to condition

v/l=3v/(4L)

1/l=3/(4*5)

l=20/3m

l=6.66m